OpenStudy (anonymous):

Swimming coach - Weekly pay: $143.75; What is her annual salary?

OpenStudy (anonymous):

143.75 x 4 x 12

OpenStudy (cccmpt2):

143.75 * 52 (52 weeks per year) = $7,475.00

OpenStudy (anonymous):

wow 2 diffrent numbers :/

OpenStudy (cccmpt2):

there are 52 weeks in a year, and there are 12 months in a year. but there aren't always only 4 weeks in a month - 4 * 12 = 48.

OpenStudy (anonymous):

My bad

OpenStudy (cccmpt2):

no prob ;)

OpenStudy (anonymous):

oo Ii c :)

OpenStudy (anonymous):

w abt.. Tour guide - Annual salary: $25,304.50; Gross pay: $973.25; How long is the pay period?

OpenStudy (anonymous):

26 times a year, which means you get paid twice a month, because there are 52 weeks in a year.

OpenStudy (cccmpt2):

25,304.50 ÷ 973.25 = 26 weeks of paychecks ÷ 52 weeks in a year = .5, or half a year. so, paychecks are given every 2 weeks = a total of 26 paychecks at 973.25 each = 25,304.50 per year.

OpenStudy (anonymous):

haha were both right

OpenStudy (anonymous):

but my answers are Bi-weekly Weekly Monthly

OpenStudy (cccmpt2):

i would say bi-weekly then, meaning 2*week...?

OpenStudy (anonymous):

26 weeks, because 2 times 26 is 56.

OpenStudy (anonymous):

Marketing consultant - Monthly pay: $1,484.25; What is his annual salary?

OpenStudy (cccmpt2):

multiply the pay by 12 (12 months per year) ;)

OpenStudy (anonymous):

Kevin Patel is a waiter at a local restaurant. He is paid $429 semimonthly. What is his annual salary?

OpenStudy (cccmpt2):

semi would equal 2 times per month. so, multiply his pay by 2 to get his montly pay, then multiply that by 12 to get his annual pay

OpenStudy (anonymous):

wait huh i dont get it

OpenStudy (cccmpt2):

he gets paid $429, 2 times per month, so that times 2 = $858 per month. then $858 8 12 months = $10,296 per year.

OpenStudy (cccmpt2):

sorry, i was eating dinner ;)